The VFlib Project

Hirotsugu Kakugawa

Hiroshima University

            Email: h.kakugawa@computer.org

1 November, 2001

Faculty of Enginnering, Hiroshima Univ. Photo by H. Kakugawa

VFlib 3.6 supports the following font file formats:

PCF, BDF, HBF, TrueType Type 1 GF, PK, Virtual Font. TFM, OFM, eKanji, Syotai Club, and JG.

VFlib Version 3.6 can be used as a font module for drivers and previewers of DVI files (TEX and L^{A}TEX). The TeX-Guy package contains a DVI interpreter (about 7300 lines of C code) and a sample DVI previewer spwx11 for X Window System written with only 140 lines of C code.

Photo by H. Kakugawa

2